# Structured Language Generation Model for Robust Structure Prediction

### 14 Feb 2024 v1

@@ Abstract
We propose Structured Language Generation Model (SLGM), a mixture of new loss function and inference method for better generalization of structured outputs. Previous studies on structure prediction (e.g. NER, RE) make use of explicit dataset information, which would boost performance, yet it might pose challenges to robust generalization in real-world situations. Instead, our model gives generalized format information about data indirectly. With format information, we could reduce sequence-to-sequence problem into classification problem via loss calibration and formatted decoding. Our experimental results showed SLGM successfully maintain performance without dataset information, and showed much less format errors. We also showed our model can work like adapters on individual dataset, with no additional training.
@@ 1 Introduction
Recent advances in pre-trained language models (PLMs) and large language models (LLMs) allow computational approaches to more tasks and problems than ever. Although they are not yet capable of human-like language learning and understanding, their general purpose and end-to-end capabilities remain an attractive option in solving natural language understanding (NLU) tasks.
@@
Previous works on structure prediction like TANL (Paolini et al., 2021) and Deepstruct reported performance improvements in various structure-related tasks with single model. We believe the improvements represent dataset-specific engineering, however, rather than improvement in structure understanding, as the improved models are fine-tuned or given explicit dataset information. Performing inference without explicit dataset information supports this, as we observe a significant drop in benchmark scores and witness a diverse array of formatting errors, as shown in Table  1 and Section  4. We aim to measure and optimize on general structure understanding abilities, rather than task-specific engineering abilities. We also focus on robustness, where the model performs consistently well in a diverse set of situations, including low-resource settings.
@@
In this paper, we propose a new framework named Structured Language Generation Model (SLGM) which can maintain robust structural understanding ability. Our framework comprise of task-specific format information, format loss, and formatted inference. The SLGM framework is illustrated in Figure 1. Task-specific information contains structure of desired outputs with tagset. Format loss works to learn the intended format, and formatted decoding encourages the model to decode with format guidelines.
We trained and evaluated our model on structure prediction tasks, demonstrating significant performance improvements over models of the same size that were trained on base sequence-to-sequence task which utilizes only cross entropy. Also, our model successfully replaced dataset information. Despite our framework not undergoing actual training, it follows performance of dataset-specific engineered models.
@@
  • We introduce SLGM, which maintain structural understanding ability well without dataset information. (Section  3)
@@ 4.3 Evaluation Method
We report score and format errors on each dataset. On every task except dialogue state tracking, we evaluated with micro F1 score. For joint entity and relation extraction, we recorded F1 score of entity and relation respectively. For dialogue state tracking task, we used joint accuracy score. We categorized format errors into three types: length mismatch, source mismatch, and tagset mismatch. Length mismatch means length of generated tuple does not match with given format. Source mismatch means the element of output violated <SOURCE> slot token restriction. When the model generated tokens that don’t exist in source sentence, source mismatch count increases. Finally, a tagset mismatch occurs when the model generates an output that is not defined within the tags specified by the dataset.
@@
NER CoNLL-03 64.03 5843 10.78 9832 84.18 25 89.79 15
OntoNotes 71.89 717 24.88 4525 81.64 99 83.84 83
GENIA 0.00 1137 66.42 26 72.24 7 67.37 35
RE TACREV 54.37 5 67.99 10 72.93 2 66.72 7
TACRED 49.18 5 60.32 10 64.32 2 57.70 7
JER CoNLL-04 0.00 794 0.00 768 71.17 0 74.43 11
4.87 530 34.24 442 33.90 2 46.43 99
NYT 85.35 475 78.60 214 89.80 95 88.83 37
54.79 82 40.85 2346 64.75 11 66.01 38
SRL CoNLL-12 80.22 20 82.31 18 84.72 0 82.15 16
ID ATIS 92.64 31 93.62 13 96.52 0 94.36 10
SNIPS 95.22 1 96.01 0 96.72 0 96.86 0
DST MultiWOZ 37.75 10 35.03 42 40.96 0 39.33 4
Average 53.10 742.31 53.16 1403.54 73.37 18.69 73.37 27.85
@@ 5.1 Comparing with plain extraction
Table  3 shows our full result. Our model showed better result than models without dataset information, and showed comparative result with dataset-aware model. Also, our model generated less format errors than dataset-aware model, which mean our model is much robust on real world dataset. When training without dataset information, we observe an average performance drop of 10 points compared to when dataset information is included. Furthermore, the frequency of format errors increases almost tenfold. Training with dataset information and then removing it yields worse results compared to training without the dataset information from the start. This means model mostly depends on dataset information.
@@ 5.2 Comparing with models using formatted decoding
To investigate the impact of formatted decoding on generating the correct format, we applied formatted decoding stage to baseline models and, conversely, removed from our SLGM model before proceeding with inference. Table  4 shows the average score and average format errors across all datasets. Formatted decoding itself can work greatly on extracting correct tagsets. On CE model trained without dataset information, model using formatted decoding achieved 16 point higher F1 score and 5% format errors compared with original model. However, when trained with dataset information, the model cannot recover from absence of dataset information even with format information. F1 score and format errors are improved, but still showed large number of format errors comparing with other models using formatted decoding. When designing models with real-world applications in mind, it becomes evident that including dataset names during training can lead to undesirable effects.
When using format loss, on the other hand, showed best score and had least format errors among models which did not use formatted decoding. We think this means format loss itself, works as dataset indicator. Slot losses produces tagset-specific loss, and this applies stronger back-propagation than cross entropy, which applies on every tokens. This makes model know more about what to extract on given sentence, which indicates the model understands structure of sentence better.
Full result and detailed analysis on each dataset is shown on Appendix  C.1
@@ 6 Additional Studies
We conducted additional studies to see in what situation our model works well. For each subsection, we report average result across all datasets. Full result is shown on Appendix  C.
@@ 6.2 Model size
When using our best methods, we achieved state of the art performance in some datasets (Table  C.3). We achieved 94.8 micro F1 score on CoNLL-03 NER dataset, and 98.3 micro F1 score on ATIS intent detection dataset, which are both state of the art. This indicates that our method can further improve through fine-tuning, and it can achieve better performance without the need for engineering specific to dataset features.
@@ 7 Conclusion and Discussion
Limitations
Format loss and formatted decoding represent basic strategies designed to compel the model to produce outputs in accordance with a predetermined format. There may be many improvements on both scoring and real-world usage. There may be many problems in specific situations. Suppose we extract named entity with location but there is no location tag in format. The model expected location and extracted some words, but the location tag is illegal. In plain decoding strategy, the model would generate any token similar to location. However, when using formatted decoding, it would generate unexpected string, which could be worse than extracting wrong tag.
@@
Ethics Statement
In our research, we utilized a pretrained model and made use of publicly available datasets published on the web. We acknowledge that the data obtained from the web may contain potential biases. We must clarify that our research did not specifically address or consider these biases. We ensured that all datasets employed in our study were accessed and used in a manner that respects their intended use and complies with any associated licenses or terms of service. We are also mindful of the potential biases present in these datasets and the pretrained model, and have endeavored to address these concerns in our research methodology.
@@ 4.2M Location
Appendix C Detailed experiment result
@@ C.1 Formatted decoding
NER CoNLL-03 64.03 5843 80.36 43 72.76 4437 84.18 25 26.80 2509 89.74 0
OntoNotes 71.89 717 78.20 191 81.30 411 81.64 99 25.36 4566 83.74 4
GENIA 0.00 1137 66.19 15 72.37 31 72.24 7 66.29 0 67.18 0
RE TACREV 54.37 5 65.21 2 72.97 7 72.93 2 67.98 3 66.74 2
TACRED 49.18 5 56.92 2 64.35 7 64.32 2 60.32 3 57.74 3
JER CoNLL-04 0.00 794 67.57 0 6.88 874 71.17 0 66.67 0 74.43 0
4.87 530 27.28 0 23.32 425 33.90 2 27.71 2 43.86 1
NYT 85.35 475 88.26 232 89.79 215 89.80 95 78.19 3 88.73 1
54.79 82 66.74 1 65.01 120 64.75 11 41.26 40 66.07 5
SRL CoNLL-12 80.22 20 82.51 1 84.72 14 84.72 0 82.30 1 82.15 1
ID ATIS 92.64 31 93.72 0 96.51 8 96.52 0 93.39 0 94.62 0
SNIPS 95.22 1 96.43 0 96.72 0 96.72 0 96.01 0 96.86 0
DST MultiWOZ 37.75 10 38.89 0 40.96 0 40.96 0 35.16 0 37.86 0
Average 53.10 742.31 69.87 37.46 66.74 503.77 73.37 18.69 59.03 548.23 73.05 1.31
@@
Table  9 shows full result using formatted decoding (Section  5.2). Comparing with Table  3, formatted decoding greatly reduces format errors on NER and JER tasks. It also shows great score improvements on low resource data like CoNLL-04.
C.2 Fine-tuning
@@
Table  10 shows full result of fine-tuning (Section  6.1). Comparing with Table  9, fine-tuning shows 5 to 7 points better F1 score than formatted decoding. Surprisingly, best setting utilizing fine-tuning is using only format loss, which is better than using both formatted decoding. Model with format loss win over model with dataset information on almost every dataset. This is evidence that our format loss can reduce problem as easy as classification. However, using both fine-tuning and formatted decoding seems making conflict to each other. This is because some gold instances violates format restriction (e.g. <SOURCE> restriction in named entity recognition). Without formatted decoding, model can infer tokens that did not appeared in source sentence, when training example exists. However, we think this is not good for real-world usage, because it can be seen as overfitting on erroneous training examples.
C.3 Model size
@@
Table  C.3 shows full result regarding to model size (Section  6.2). On cross entropy model, model can distinguish characteristics of dataset when model parameters increase. The power of formatted inference on score is reduced when parameter size increased, yet still effective reducing format errors. Comparing CE+FL with CE, using only format loss showed worse score and more format error on large sized model. However, when mixed with formatted decoding, it showed additional increase comparing with models using single methods. Additionally, when we use every method we tried in this paper (i.e. format loss, formatted decoding, using bigger model, fine-tuning on dataset), we achieved state of the art performance on CoNLL-03 named entity recognition task and ATIS intent detection task. This means our framework still has room for improvement.

### 19 Feb 2024 v2

@@ Abstract
Previous work in structured prediction (e.g. NER, information extraction) using single model make use of explicit dataset information, which helps boost in-distribution performance but is orthogonal to robust generalization in real-world situations. To overcome this limitation, we propose the Structured Language Generation Model (SLGM), a framework that reduces sequence-to-sequence problems to classification problems via methodologies in loss calibration and decoding method. Our experimental results show that SLGM is able to maintain performance without explicit dataset information, follow and potentially replace dataset-specific fine-tuning.
@@ 1 Introduction
Recent advances in pre-trained language models (PLMs) and large language models (LLMs) allow computational approaches to more tasks and problems than ever. LLMs have demonstrated exceptional end-to-end problem-solving capabilities; however, their internal workings remain largely a black box. Therefore, tasks related to natural language understanding (NLU) still remain significant from the perspective of explainable AI  (Ribeiro et al., 2016).
@@
Previous works on structure prediction like TANL (Paolini et al., 2021) and Deepstruct reported performance improvements in various structure-related tasks with a single model. However, a nontrivial portion of such improvements represent successful dataset-specific engineering, rather than an overall improvement in structure understanding, as the improved models rely on fine-tuning and explicit dataset information. Performing inference without explicit dataset information supports this, as we observe a significant drop in benchmark scores and witness a diverse array of formatting errors, as shown in Table  1 and Section  4. We aim to measure and optimize on general structure understanding abilities, rather than task-specific engineering abilities. We also focus on robustness, where the model performs consistently well in a diverse set of situations, including low-resource settings.
@@
In this paper, we propose a new framework named Structured Language Generation Model (SLGM) which can maintain robust structural understanding ability. Our framework comprises of task-specific format information, format loss, and formatted inference. The SLGM framework is illustrated in Figure 1. Task-specific information contains structure of desired output structure, which constitutes the task-specific restrictions along with possible labels. Format loss works to learn the intended format by penalizing incorrect output format. Finally, formatted decoding encourages the model to decode with format guidelines.
We train and evaluate SLGM on structure prediction tasks. Our experimental results demonstrates significant performance improvements over baseline models of same size which optimizes only with cross entropy loss. Also, SLGM follows performance of models with explicit dataset information and models fine-tuned on specific dataset, which indicates our method can replace dataset-specific engineering.
@@
  • We introduce SLGM, which is able to maintain structural understanding ability well without dataset information. (Section  3)
@@ 4.3 Evaluation Method
We report score and format errors on each dataset. On every task except dialogue state tracking, we evaluated with micro F1 score. For joint entity and relation extraction, we recorded F1 score of entity and relation respectively. For dialogue state tracking task, we used joint accuracy score. We categorized format errors into three types: length mismatch, source mismatch, and tagset mismatch. Length mismatch means length of generated tuple does not match with given format. Source mismatch means the element of output violated <SOURCE> slot token restriction. When the model generated tokens that don’t exist in source sentence, source mismatch count increases. Finally, a tagset mismatch occurs when the model generates an output that is not defined within the tags specified by the dataset. We report results averaged on 5 runs.
@@
NER CoNLL-03 69.32 4700 12.06 10273 80.28 43 87.11 5
OntoNotes 75.44 766 24.79 5143 75.87 142 81.12 348
GENIA 67.05 147 65.26 70 69.88 5 66.48 30
RE TACREV 66.95 6 66.52 21 71.39 2 67.04 13
TACRED 58.41 6 59.51 21 63.11 2 59.07 13
JER CoNLL-04 0.00 873 0.00 827 71.74 0 74.88 14
13.31 413 23.08 491 27.87 2 48.53 46
NYT 88.68 322 74.60 412 88.80 149 88.45 23
65.80 17 45.07 464 59.36 20 67.47 7
SRL CoNLL-12 82.47 161 82.37 158 83.45 0 82.35 159
ID ATIS 94.09 11 94.30 15 93.96 0 94.21 9
SNIPS 96.58 0 95.79 1 96.86 0 96.07 1
DST MultiWOZ 38.16 667 36.68 0 38.87 0 37.18 0
Average 62.79 622.23 52.31 1376.62 70.88 28.08 73.07 51.38
@@ 5.1 Comparing with plain extraction
Table  3 shows our full result. Our model showed better result than models without dataset information, and showed comparative result with dataset-aware model on most dataset. Also, our model generated less format errors than dataset-aware model, which mean our model is much robust on real world dataset. When training without dataset information, we observe an average performance drop of 11 points compared to when dataset information is included. Furthermore, the frequency of format errors increases 12 times larger. Training with dataset information and then removing it yields worse results compared to training without the dataset information from the start. This means model mostly depends on dataset information.
@@ 5.2 Comparing with models using formatted decoding
To investigate the impact of formatted decoding on generating the correct format, we applied formatted decoding stage to baseline models and, conversely, removed from our SLGM model before proceeding with inference. Table  4 shows the average score and average format errors across all datasets. Formatted decoding itself can work greatly on extracting correct tagsets. On CE model trained without dataset information, model using formatted decoding achieved 6 point higher F1 score and 7% format errors compared with original model. However, when trained with dataset information, the model cannot recover from absence of dataset information even with format information. F1 score and format errors are improved, but still showed large number of format errors comparing with other models using formatted decoding. When designing models with real-world applications in mind, it becomes evident that including dataset names during training can lead to undesirable effects.
When comparing models without formatted decoding, model with format loss showed best score and least format errors. We think this means format loss itself, works as dataset indicator. Slot losses produces tagset-specific loss, and this applies stronger back-propagation than cross entropy, which applies on every tokens. This makes model know more about what to extract on given sentence, which indicates the model understands structure of sentence better.
Full result on each dataset is shown on Appendix  D.1
@@ 6 Additional Studies
We conducted additional studies to see in what situation our model works well. In this section, we report average result across all datasets with single run. Full result is shown on Appendix  D.
@@ 6.2 Model size
When using our best methods, we achieved state of the art performance in some datasets (Table  D.3). We achieved 94.8 micro F1 score on CoNLL-03 NER dataset, and 98.3 micro F1 score on ATIS intent detection dataset, which are both state of the art. This indicates that our method can further improve through fine-tuning, and it can achieve better performance without the need for engineering specific to dataset features.
@@ 7 Conclusion and Discussion
8 Limitations
Format loss and formatted decoding represent basic strategies designed to compel the model to produce outputs in accordance with a predetermined format. There may be many improvements on both scoring and real-world usage. There may be many problems in specific situations. Suppose we extract named entity with location but there is no location tag in format. The model expected location and extracted some words, but the location tag is illegal. In plain decoding strategy, the model would generate any token similar to location. However, when using formatted decoding, it would generate unexpected string, which could be worse than extracting wrong tag. We did not conduct quantitative experiments regarding this situation.
@@ 8 Limitations
9 Ethics Statement
In our research, we utilized a pretrained model and made use of publicly available datasets published on the web. We acknowledge that the data obtained from the web may contain potential biases. Our model can bear potential risks and harms discussed in  (Brown et al., 2020), and we clarify that our research did not specifically address or consider these risks. We ensured that all datasets employed in our study were accessed and used in a manner that respects their intended use and complies with any associated licenses or terms of service. We are also mindful of the potential biases present in these datasets and the pretrained model.
@@ 4.2M Location
Appendix C Dataset Statistics
Task Dataset Train Dev Test
NER CoNLL-03 14986 3465 7148
Ontonotes-v5 75187 9603 9479
GENIA 15023 1669 1854
JER CoNLL-04 922 231 288
NYT 56196 5000 5000
RE TACRED* 68124 22631 15509
SRL CoNLL-12 253070 35297 24462
ID ATIS 4478 500 893
SNIPS 13084 700 700
DST MultiWOZ 62367 7371 7368
Table 9: Statistics of each multi-task dataset. Each cell represents sentence number.
Table  9 shows statistics of multi-task dataset. Note that TACRED and TACREV shares same number of sentences; their difference is on validation and test labels. On training with format loss, we removed sentences which have labels violating format errors (less than 1%). On evaluation, we used every sentences as-is.
Appendix D Detailed experiment result
@@ D.1 Formatted decoding
NER CoNLL-03 69.32 4700 78.09 79 66.17 5127 80.28 43 28.30 2196 87.11 0
OntoNotes 75.44 766 76.04 288 75.50 572 75.87 142 26.72 5164 81.64 45
GENIA 67.05 147 66.31 9 70.31 75 69.88 5 64.95 3 66.35 1
RE TACREV 66.95 6 66.93 2 71.51 12 71.39 2 66.48 2 67.00 2
TACRED 58.41 6 58.42 2 63.21 12 63.11 2 59.48 2 59.05 2
JER CoNLL-04 0.00 873 66.73 0 0.00 873 71.74 0 66.02 0 74.50 0
13.31 413 26.49 3 16.84 493 27.87 2 27.23 2 45.71 2
NYT 88.68 322 88.45 156 88.84 300 88.80 149 73.79 17 88.36 1
65.80 17 65.75 0 59.42 33 59.36 20 45.03 8 67.47 2
SRL CoNLL-12 82.47 161 82.35 3 83.44 19 83.45 0 82.27 1 82.24 1
ID ATIS 94.09 11 93.96 0 94.14 12 93.96 0 93.85 0 94.07 0
SNIPS 96.58 0 96.58 0 96.86 0 96.86 0 95.72 0 96.01 0
DST MultiWOZ 38.16 667 38.21 0 38.85 7 38.87 0 36.68 0 37.14 0
Average 62.79 622.23 69.56 41.69 63.47 579.62 70.88 28.08 58.96 568.85 72.82 4.31
@@
Table  10 shows full result using formatted decoding (Section  5.2), averaged on 5 runs. Comparing with Table  3, formatted decoding greatly reduces format errors on NER and JER tasks. It also shows great score improvements on low resource data like CoNLL-04. Even if the model knows about dataset information, formatted decoding still upgrades score and reduces format errors.
D.2 Fine-tuning
@@
Table  11 shows full result of fine-tuning (Section  6.1) on single run. Comparing with Table  10, fine-tuning shows 5 to 7 points better F1 score than formatted decoding. Surprisingly, best setting utilizing fine-tuning is using only format loss, which is better than using both formatted decoding. Model with format loss win over model with dataset information on almost every dataset. This is evidence that our format loss can reduce problem as easy as classification. However, using both fine-tuning and formatted decoding seems making conflict to each other. This is because some gold instances violates format restriction (e.g. <SOURCE> restriction in named entity recognition). Without formatted decoding, model can infer tokens that did not appeared in source sentence, when training example exists. However, we think this is not good for real-world usage, because it can be seen as overfitting on erroneous training examples.
D.3 Model size
@@
Table  D.3 shows full result regarding to model size (Section  6.2) on single run. On cross entropy model, model can distinguish characteristics of dataset when model parameters increase. The power of formatted inference on score is reduced when parameter size increased, yet still effective reducing format errors. Comparing CE+FL with CE, using only format loss showed worse score and more format error on large sized model. However, when mixed with formatted decoding, it showed additional increase comparing with models using single methods. Additionally, when we use every method we tried in this paper (i.e. format loss, formatted decoding, using bigger model, fine-tuning on dataset), we achieved state of the art performance on CoNLL-03 named entity recognition task and ATIS intent detection task. This means our framework still has room for improvement.